# Fractional binomial distributions induced by the generalized binomial theorem and their applications

### v1

@@
Abstract.
A generalized binomial theorem was established by Hara and Hino [Bull. London Math. Soc. 42 (2010), 467–477] for proving the neo-classical inequality. We introduce a new fractional analogue of the binomial distribution on the basis of the generalized binomial theorem and study its properties. They include the law of large numbers, the central limit theorem, and the law of small numbers, the last of which leads to a fractional analogue of the Poisson distribution. As an application of the fractional binomial distribution, a fractional analogue of the Bernstein operator is proposed, which acts on the space of continuous functions on the unit interval. We show that the fractional Bernstein operator uniformly approximates every continuous function on the unit interval. Some limit theorems for its iterates are discussed as well. In particular, a generalization of the Wright–Fisher diffusion process arising in the study of population genetics is captured as a certain limit of the iterates.
Key words and phrases:
fractional binomial distribution, neo-classical inequality, fractional Taylor expansion, fractional Bernstein operator
2020 Mathematics Subject Classification:
Primary 60E05; Secondary 26A33, 60F05, 41A36.
1. Introduction
1.1. The neo-classical inequality and the generalized binomial theorem
Lyons [Lyo98] initiated the theory of rough paths and introduced the neo-classical inequality in order to show a fundamental theorem of rough path theory (see [Lyo98, Theorem 2.2.1]). This asserts that
$\alpha^{2}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}y^{\alpha(n-j)}\leq(x+y)^{\alpha n}$ (1.1)
for $\alpha\in(0,1]$, $n\in\mathbb{N}$ and $x,y\geq 0$. Here, we define
@@ 1.1. The neo-classical inequality and the generalized binomial theorem
for $w\in\mathbb{C}\setminus\{-k\mid k\in\mathbb{N}\}$ and $z\in\mathbb{C}$ in general, where $\Gamma(\cdot)$ denotes the gamma function. If $\infty$ appears in the denominator of the right-hand side of the above, then $\binom{w}{z}$ is regarded as zero. Also, $0^{0}$ is defined as 1. The equality holds in (1.1) if $\alpha=1$, which is nothing but the usual binomial theorem. Lyons also conjectured that the coefficient $\alpha^{2}$ in (1.1) can be replaced by $\alpha$. Afterwards, Hara and Hino [HH10] gave an affirmative answer to Lyons’ conjecture.
Proposition 1.1 (Neo-classical inequality, [HH10, Theorem 1.2]).
Let $\alpha\in(0,1]$. For $n\in\mathbb{N}$ and $x,y\geq 0$, we have
$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}y^{\alpha(n-j)}\leq(x+y)^{\alpha n}.$ (1.2)
The equality holds if and only if $\alpha=1$ or $x=y=0$.
When $\alpha\in(0,1)$ and only one of $x$ and $y$ is zero, the inequality (1.2) is trivial with strict inequality. Therefore, in order to show (1.2), it suffices to consider the case where $\alpha\in(0,1)$ and $x,y>0$. Moreover, we may assume that $x\leq y$ by the symmetry of the roles of $x$ and $y$ in (1.2). By dividing both sides by $y^{\alpha n}$ and letting $\lambda=x/y$, (1.2) is equivalent to
$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}\leq(1+\lambda)^{\alpha n},\qquad\alpha\in(0,1),\,n\in\mathbb{N},\,\lambda\in(0,1].$ (1.3)
For proving (1.3), Hara and Hino established a generalization of the binomial theorem [HH10, Theorem 1.3], which is stated roughly as follows:
$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}\lambda^{\alpha j}=(1+\lambda)^{\alpha n}+R_{\alpha}^{(n)}(\lambda),\qquad\alpha\in(0,2),\,n\in\mathbb{N},\,\lambda\in(0,1],$ (1.4)
where the remainder term $R_{\alpha}^{(n)}(\lambda)$ is given explicitly. Indeed, by the expression, the term $R_{\alpha}^{(n)}(\lambda)$ is strictly negative when $\alpha\in(0,1)$, which readily implies (1.3).
Furthermore, they extended the generalized binomial theorem (1.4) to the case where $\alpha$ is an arbitrary positive number. Let us now give the precise statement. In general, for $z\in\mathbb{C}\setminus\{x\in\mathbb{R}\mid x\leq 0\}$ and $\gamma\in\mathbb{R}$, the power $z^{\gamma}$ is defined as $\exp(\gamma\,\mathrm{Log}\,z)$, where $\mathrm{Log}\,z$ is the principal value of $\log z$ so that Log $\mathrm{Log}\,1=0$. For $\alpha>0$, we put
We note that $K_{\alpha}=\{1\}$ when $\alpha\in(0,2)$ and that $-1\in K_{\alpha}$ if and only if $\alpha/2\in\mathbb{N}$. Then, the generalized binomial theorem is stated as follows.
Proposition 1.2 ([HH10, Theorem 3.2]).
@@
×$\hskip 28.45274pt\times\left\{\frac{1}{|t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi}|^{2}}+\frac{\lambda^{\alpha n}}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\right\}\,\mathrm{d}t.$ (1.6)
@@
It should be noted that the first term on the right-hand side of (1.2) is a real number because $\omega\in K_{\alpha}$ implies $\overline{\omega}\in K_{\alpha}$. For proving Proposition 1.2, the authors provided in [HH10, Theorem 3.1] the fractional-order Taylor-like series expansions with residual terms by using basic complex analysis. See Proposition 2.1 for further details.
1.2. Fractional binomial distribution
Suppose that $n\in\mathbb{N}$ and $x\in[0,1]$. The binomial distribution $\mu_{x}^{(n)}$ defined as
$\mu_{x}^{(n)}(\{j\})=\binom{n}{j}x^{j}(1-x)^{n-j},\qquad j=0,1,2,\dots,n$
on $\{0,1,2,\dots,n\}$ is one of the most important discrete probability distributions. The equality $\sum_{j=0}^{n}\mu_{x}^{(n)}(\{j\})=1$ is derived from the usual binomial theorem. Because Proposition 1.2 is regarded as a generalization of the binomial theorem, it is quite natural to ask for a fractional analogue of the binomial distribution in accordance with Proposition 1.2.
For $\alpha>0$, $n\in\mathbb{N}$ and $x\in(0,1)$, we have
@@ 1.2. Fractional binomial distribution
×$\hskip 28.45274pt\times\left\{\frac{(1-x)^{\alpha n}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}+\frac{x^{\alpha n}}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\right\}\,\mathrm{d}t$ (1.7)
by putting $\lambda=x/(1-x)$ in (1.2) and multiplying both sides by $(1-x)^{\alpha n}$. Keeping in mind that the right-hand side of (1.7) is close to 1 when $n$ is sufficiently large, this identity gives rise to a discrete probability distribution on $\{0,1,2,\dots,n\}$ as follows.
@@
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in(0,1)$. Then, a probability measure $\mu_{\alpha,x}^{(n)}$ on $\{0,1,2,\dots,n\}$, called the $\alpha$-fractional binomial distribution, is defined as
$\mu_{\alpha,x}^{(n)}(\{j\})=\frac{\alpha}{Z_{\alpha,x}^{(n)}}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)},\qquad j=0,1,2,\dots,n.$
Moreover, we define $Z_{\alpha,0}^{(n)}=Z_{\alpha,1}^{(n)}=\alpha$, $\mu_{\alpha,0}^{(n)}=\delta_{0}$, and $\mu_{\alpha,1}^{(n)}=\delta_{1}$, where $\delta_{x}$ is the Dirac measure at $x$.
It is clear that $\mu_{1,x}^{(n)}$ coincides with the binomial distribution $\mu_{x}^{(n)}$. The probability distribution $\mu_{\alpha,x}^{(n)}$ is regarded as a fractional variant of $\mu_{x}^{(n)}$. As seen in Corollary 2.8, one of the features is that the mean of $\mu_{\alpha,x}^{(n)}$ is asymptotically equal to that of $\mu_{x}^{(n)}$ as $n\to\infty$, while the variance of $\mu_{\alpha,x}^{(n)}$ is asymptotically $1/\alpha$ times that of $\mu_{x}^{(n)}$.
One of the aims of this study is to study some basic properties of the $\alpha$-fractional binomial distribution $\mu_{\alpha,x}^{(n)}$, and all such properties are given with their proofs in Section 2. In Section 2.1, we review the fractional-order Taylor-like series of holomorphic functions $f$ with residual terms [HH10, Theorem 2.1]. See Proposition 2.1 for its precise statement. In Section 2.2, we provide explicit expressions for moments of $\mu_{\alpha,x}^{(n)}$ by making effective use of the fractional-order Taylor-like series expansion. See Theorem 2.5 and Corollary 2.8. In particular, we see that the ($c$-)Whitney numbers of the second kind—which in combinatorics are a certain generalization of the Stirling numbers of the second kind—appear in the expression for each moment of $\mu_{\alpha,x}^{(n)}$. Section 2.3 discusses a uniform bound of the normalized constant $Z_{\alpha,x}^{(n)}$ with respect to both $n\in\mathbb{N}$ and $x\in[0,1]$. Furthermore, in Section 2.4, a uniform estimate for the moments of $\mu_{\alpha,x}^{(n)}$ is also obtained. Section 2.5 focuses on the characteristic function $\varphi_{\alpha,x}^{(n)}(\xi)$ of $\mu_{\alpha,x}^{(n)}$, which plays a crucial role in obtaining the central limit theorem for $\mu_{\alpha,x}^{(n)}$. Lastly, in Section 2.6, a multivariate version of the neo-classical inequality is deduced. Similarly to Definition 1.3, this inequality also induces a fractional analogue of the multinomial distribution. Being different from the binomial case, however, it seems quite difficult to obtain even basic characteristics of the fractional multinomial distribution such as its mean vector and its variance-covariance matrix, because the multivariate version of the fractional Taylor series expansion is yet to be obtained. The discussion of the multivariate cases is left to the future.
Here, note that the terminology fractional in probability theory is used in several different contexts. For instance, a typical use can be seen in the notion of fractional Brownian motion. See [Lee21] (for example) for the fractional binomial distribution in this sense. Yet another use is derived from fractional calculus. We refer to [CP12] for the fractional binomial process defined in terms of Caputo’s fractional derivatives. However, to the best of our knowledge, no paper to date has discussed the fractional binomial distribution in our sense, which means that the present paper gives a new insight into the study of probability theory and other areas of mathematics.
Based on the properties given in Section 2, we give analogues of three classical limit theorems in Section 3. Sections 3.1 and 3.2 are devoted to the proofs of the weak law of large numbers and the central limit theorem for the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$. We also obtain the law of small numbers for $\mu_{\alpha,x}^{(n)}$ in Section 3.3. Namely, we show that $\mu_{\alpha_{n},x_{n}}^{(n)}$ converges weakly to a certain probability distribution with masses on $\{0,1,2,\dots\}$ as $n\to\infty$ together with $\alpha_{n}\to\alpha$ and $\alpha_{n}nx_{n}\to\lambda$ for some $\alpha>0$ and $\lambda>0$. The limiting probability distribution $\nu_{\alpha,\lambda}$ is regarded as a fractional analogue of the Poisson distribution. See Definition 3.3 for the precise definition. Obviously, $\mu_{1,\lambda}$ coincides with the usual Poisson distribution. Note that $\nu_{\alpha,\lambda}$ is defined in terms of a special function called the Mittag-Leffler function $E_{\alpha}(z)$, $\alpha>0$, which is a generalization of the exponential function. The distribution $\nu_{\alpha,\lambda}$ is a special case of generalized fractional Poisson distributions discussed in [Her16, CO17]. We also mention that the Mittag-Leffler function has already been used to introduce some probability distributions in different situations. We refer to [Pil90, PJ95] (for example) for the notion of the Mittag-Leffler distribution, although it is far from the Poisson distribution. A few basic properties of the fractional Poisson distribution such as a recurrence formula for its moments are discussed in Section 3.4 as well. Because the Mittag-Leffler function is used frequently in fractional calculus, further investigations of $\nu_{\alpha,\lambda}$ from the point of view of fractional calculus are expected.
1.3. Fractional Bernstein operator
There exist various applications of the binomial distribution to other branches of mathematics. One of them is seen in the study of positive linear operators acting on $C([0,1])$, the Banach space of all continuous functions on $[0,1]$ endowed with the supremum norm $\|\cdot\|_{\infty}$. Bernstein introduced in his short paper [Ber12] the following positive linear operator acting on $C([0,1])$ defined in terms of the binomial distribution.
Definition 1.4 (Bernstein operator, cf. [Ber12]).
For $n\in\mathbb{N}$, a positive linear operator $B_{n}$ acting on $C([0,1])$ is defined by
$B_{n}f(x)=\sum^{n}_{k=0}\binom{n}{k}x^{k}(1-x)^{n-k}f\left(\frac{k}{n}\right),\qquad f\in C([0,1]),\,x\in[0,1].$ (1.8)
The positive linear operator $B_{n}$ is called the Bernstein operator.
The right-hand side of (1.8) is called the Bernstein polynomial. Bernstein proved in [Ber12] that $B_{n}$ uniformly approximates every continuous function $f\in C([0,1])$.
Proposition 1.5 (cf. [Ber12]).
For any $f\in C([0,1])$, we have
$\lim_{n\to\infty}\|B_{n}f-f\|_{\infty}=0.$
It is clear that Proposition 1.5 gives a constructive proof of the celebrated Weierstrass approximation theorem.
It is also known that the Bernstein operator has a number of properties of interest. We refer to [Bus17] for comprehensive information on the Bernstein operator with extensive references therein, and to [Alt10, AC94] for surveys on approximating properties of positive linear operators acting on some function spaces, including the Bernstein operator. On the other hand, it is natural to consider limiting behaviors of $B_{n}^{k}$, the $k$-times iterations of $B_{n}$, as $k\to\infty$. One such limiting behavior was first investigated in [KR67].
Proposition 1.6 (cf. [KR67, Theorem 1]).
Let $n\in\mathbb{N}$ be fixed. Then, we have
$\lim_{k\to\infty}\max_{x\in[0,1]}|B_{n}^{k}f(x)-\{f(1)x+f(0)(1-x)\}|=0,\qquad f\in C([0,1]).$
Since the publication of [KR67], several authors have tried to make the proof of Proposition 1.6 more sophisticated. We refer to [AI09, Rus04] (for example) for its short proofs. Moreover, Konstantopoulos, Yuan and Zazanis [KYZ18] discussed a limit theorem for iterates of $B_{n}$ with scales different from Proposition 1.6. In particular, they gave a probabilistic proof of the following limit theorem.
Proposition 1.7 (cf. [KYZ18, Theorem 3]).
For $f\in C([0,1])$ and $t\geq 0$, we have
$\lim_{n\to\infty}\|B_{n}^{\lfloor nt\rfloor}f-\mathbb{E}[f(X_{\bullet}(t))]\|_{\infty}=0,$
where $(X_{x}(t))_{t\geq 0}$ is the Wright–Fisher diffusion process which solves
$\mathrm{d}X_{x}(t)=\sqrt{X_{x}(t)(1-X_{x}(t))}\,\mathrm{d}W(t),\qquad X_{x}(0)=x\in[0,1].$
Here, $(W(t))_{t\geq 0}$ is standard one-dimensional Brownian motion.
This theorem asserts that the sequence of discrete semigroups $(B_{n}^{\lfloor nt\rfloor})_{t\geq 0}$, $n\in\mathbb{N}$ converges to the Wright–Fisher diffusion semigroup as $n\to\infty$. In other words, the properly scaled Markov chain $X_{x}^{(n)}(\lfloor nt\rfloor)$ whose one-step transition operator is given by $B_{n}$ converges in law to $X_{x}(t)$ as $n\to\infty$ for each $t\geq 0$. We also refer to [KZ70] for early works on limit theorems for iterates of some positive linear operators. For further details on the Wright–Fisher diffusion, see [EK86, Section 10] (for example).
In Section 4, we introduce a fractional analogue of the Bernstein operator with the binomial distribution replaced by the fractional one defined in Definition 1.3, and we investigate some of its properties including limit theorems for its iterates. The following is the definition of the fractional Bernstein operator.
@@ 1.3. Fractional Bernstein operator
Let $\alpha>0$ and $n\in\mathbb{N}$. The $\alpha$-fractional Bernstein operator $B_{\alpha,n}$ is defined by
$B_{\alpha,n}f(x)$ $:=\sum_{j=0}^{n}\mu_{\alpha,x}^{(n)}(\{j\})f\left(\frac{j}{n}\right)$
$=\sum_{j=0}^{n}\frac{\alpha}{Z_{\alpha,x}^{(n)}}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}f\left(\frac{j}{n}\right),\qquad f\in C([0,1]),\,x\in[0,1].$
We note that $B_{1,n}$ coincides with the Bernstein operator $B_{n}$.
In Section 4.1, we show in Theorem 4.1 that the fractional Bernstein operator $B_{\alpha,n}$ also uniformly approximates every continuous function in $C([0,1])$ similarly to Proposition 1.5. Moreover, Sections 4.2 and 4.3 discuss limiting behaviors of the iterates $B_{\alpha,n}^{k}f$, $f\in C([0,1])$ as $k\to\infty$. In the former section, we obtain the Kelisky–Rivlin-type limit theorem for $B_{\alpha,n}$. In the latter section, we concentrate on a fractional analogue of Proposition 1.7. We first show that the sequence $\{n(B_{\alpha,n}-I)\}_{n=1}^{\infty}$ converges to a second-order differential operator $\mathcal{L}_{\alpha}=(2\alpha)^{-1}x(1-x)f^{\prime\prime}(x)$ on $C^{2}([0,1])$ with Neumann boundary condition. This means the convergence of the infinitesimal generator of the discrete Markov semigroup generated by $B_{\alpha,n}$ to that of a diffusion semigroup. Note that $\mathcal{L}_{\alpha}$ is the generator of the Wright–Fisher diffusion with diffusion coefficient different from that of $(X_{x}(t))_{t\geq 0}$ up to a multiple constant. Next, we show the uniform convergence of the iterates $B_{\alpha,n}^{\lfloor nt\rfloor}f$, $f\in C([0,1])$ to the Wright–Fisher semigroup corresponding to $\mathcal{L}_{\alpha}$ as $n\to\infty$, by employing the celebrated Trotter approximation theorem (cf. [Tro58]). See Lemma 4.5 and Theorem 4.4 for precise statements. We also discuss in Section 4.4 that Theorem 4.4 holds even though we replace the parameter $\alpha>0$ with a continuous function $\alpha(x)\colon[0,1]\to(0,\infty)$. In this case, the corresponding infinitesimal generator is given by $\mathcal{L}_{\alpha(\cdot)}=(2\alpha(x))^{-1}x(1-x)f^{\prime\prime}(x)$ on $C^{2}([0,1])$. Hence, it turns out that the stochastic processes captured by the limit of $B_{\alpha(\cdot),n}^{\lfloor nt\rfloor}$ as $n\to\infty$ are not only Wright–Fisher diffusions but also more-general diffusions on $[0,1]$.
Here, we mention some existing notions of linear operators named the fractional Bernstein operator. In [Man15], the author introduced the fractional Bernstein operator $(1-B_{n})^{\alpha}$, $\alpha>0$ in order to apply density estimations of probability distributions in statistics and data science, which differs completely from ours. Another different formulation of the fractional Bernstein operator from ours can be found in [Yu21]. Our fractional Bernstein operator $B_{\alpha,n}$ seems new in that we define it via the generalized binomial theorem deriving the neo-classical inequality. We expect to develop further studies of $B_{\alpha,n}$ in the near future.
1.4. Notation

### v3

@@
Abstract
We develop a fractional extension of the classical binomial distribution and the associated Bernstein operator, formulated within the framework of the generalized binomial theorem (Hara and Hino [Bull. London Math. Soc. 42 (2010), 467–477]). This provides a new probabilistic structure not representable as the law of the sum of independent and identically distributed random variables. Despite this nonstandard nature, we establish several of its fundamental analytic and probabilistic properties, including limit theorems, through a unified framework based on the generalized binomial theorem. We further analyze the properties of the fractional Bernstein operator associated with the fractional binomial distribution. In particular, we prove that the iterates of the operator converge to a generalized Wright–Fisher diffusion semigroup after a proper diffusive rescaling.
keywords:
fractional binomial distribution, fractional Taylor expansion, fractional Bernstein operator
pacs:
MSC Classification]60E05, 26A33, 60F05, 41A36.
1 Introduction
1.1 Fractional binomial distribution and fractional Bernstein operator
The binomial distribution and the associated Bernstein operator form one of the classical bridges between probability and approximation theory. The classical binomial distribution
$\mu_{x}^{(n)}(\mathrm{d}z):=\sum_{j=0}^{n}\binom{n}{j}x^{j}(1-x)^{n-j}\delta_{j}(\mathrm{d}z),\qquad n\in\mathbb{N},\ x\in(0,1),$
arises naturally from the classical binomial theorem $(x+y)^{n}=\sum_{j=0}^{n}\binom{n}{j}x^{j}y^{n-j}$. Here, $\delta_{j}$ denotes the Dirac measure at $j$. It corresponds to the law of the sum of independent and identically distributed (i.i.d. in short) Bernoulli random variables, and induces the celebrated Bernstein operator, acting on the set $C([0,1])$ consisting of continuous functions on $[0,1]$, defined as
$B_{n}f(x):=\int_{\mathbb{R}}f\left(\frac{z}{n}\right)\,\mu_{x}^{(n)}(\mathrm{d}z)=\sum_{j=0}^{n}\binom{n}{j}x^{j}(1-x)^{n-j}f\left(\frac{j}{n}\right),\qquad f\in C([0,1]).$
Bernstein proved in [3] that $\|B_{n}f-f\|_{\infty}\to 0$ holds as $n\to\infty$ for $f\in C([0,1])$, where $\|\cdot\|_{\infty}$ denotes the supremum norm. We refer to e.g., [4, 1, 2] for further details. Moreover,
$B_{n}^{k}:=\underbrace{B_{n}\circ\cdots\circ B_{n}}_{k},\qquad k\in\mathbb{N},$
forms a discrete semigroup, whose certain scaling limit as $n\to\infty$ yields the Wright–Fisher diffusion semigroup studied in population genetics. See [15, 7, 18].
While such a classical theory has been studied intensively and extensively in probability theory, the fractional analogues of such structures remain largely unexplored. It is worth noting that the generalized binomial theorem obtained by Hara–Hino [12] provides a natural framework for such an extension. In general, for $z\in\mathbb{C}\setminus\{x\in\mathbb{R}\mid x\leq 0\}$ and $\gamma\in\mathbb{R}$, the power $z^{\gamma}$ is defined as $\exp(\gamma\,\mathrm{Log}\,z)$, where $\mathrm{Log}\,z$ is the principal value of $\log z$ so that $\mathrm{Log}\,1=0$. For $\alpha>0$, put
Note that $K_{\alpha}=\{1\}$ when $\alpha\in(0,2)$ and that $1\in K_{\alpha}$ if and only if $\alpha/2\in\mathbb{N}$. We define
@@ 1.1 Fractional binomial distribution and fractional Bernstein operator
for $w\in\mathbb{C}\setminus\{-k\mid k\in\mathbb{N}\}$ and $z\in\mathbb{C}$ in general, where $\Gamma(\cdot)$ denotes the gamma function. If $\infty$ appears in the denominator of the right-hand side of the above, then $\binom{w}{z}$ is regarded as zero. Also, $0^{0}$ is defined as 1. The generalized binomial theorem is stated as follows.
Proposition 1.1 ([12, Theorem 3.2]).
@@
$\quad\times\left\{\frac{1}{|t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi}|^{2}}+\frac{\lambda^{\alpha n}}{|e^{-i\alpha\pi}-(\lambda t)^{\alpha}|^{2}}\right\}\,\mathrm{d}t.$ (1.2)
@@
We note that Proposition 1.1 also holds for $\lambda>0$ by analytic continuation in the parameter $\lambda$. For proving Proposition 1.1, the authors established the fractional-order Taylor-like series expansions with residual terms in [12, Theorem 3.1], which has the fractional calculus as its theoretical background. See Proposition 2.1 for further details. Incidentally, Proposition 1.1 is used to answer Lyons’ conjecture relating to the neo-classical inequality, which plays a crucial role in the proof of a fundamental theorem of rough path theory (see [22, Theorem 2.2.1]).
Let $\alpha>0$ and $n\in\mathbb{N}$. By putting $\lambda=x/(1-x)$, $x\in(0,1)$, in (1.1) and multiplying both sides by $(1-x)^{\alpha n}$, we have
@@
$\quad\times\left\{\frac{(1-x)^{\alpha n}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}+\frac{x^{\alpha n}}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\right\}\,\mathrm{d}t.$ (1.3)
Keeping in mind that the right-hand side of (1.3) is close to 1 when $n$ is sufficiently large, we undertake a systematic study of the fractional binomial distribution and its associated fractional Bernstein operator, which are defined as follows.
@@
Let $\alpha>0$, $n\in\mathbb{N}$, and $x\in[0,1]$. Then, a probability measure $\mu_{\alpha,x}^{(n)}$ defined as
$\mu_{\alpha,x}^{(n)}(\mathrm{d}z):=\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\delta_{j}(\mathrm{d}z)$
is called the $\alpha$-fractional binomial distribution.
Note that $Z_{\alpha,0}^{(n)}=Z_{\alpha,1}^{(n)}=\alpha$, $\mu_{\alpha,0}^{(n)}=\delta_{0}$, and $\mu_{\alpha,1}^{(n)}=\delta_{1}$.
@@
Let $\alpha>0$ and $n\in\mathbb{N}$. The $\alpha$-fractional Bernstein operator $B_{\alpha,n}$, acting on $C([0,1])$, is defined by
$B_{\alpha,n}f(x):\!$ $=\int_{\mathbb{R}}f\left(\frac{z}{n}\right)\,\mu_{\alpha,x}^{(n)}(\mathrm{d}z)$
$=\frac{\alpha}{Z_{\alpha,x}^{(n)}}\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}f\left(\frac{j}{n}\right),\qquad f\in C([0,1]),\,x\in[0,1].$
It is clear that $\mu_{1,x}^{(n)}$ and $B_{1,n}$ coincide with $\mu_{x}^{(n)}$ and $B_{n}$, respectively. In this sense, the fractional binomial distribution (resp. fractional Bernstein operator) is a natural generalization of the binomial distribution (resp. Bernstein operator) with additional parameter $\alpha$. This family interpolates smoothly between the classical and the newly introduced fractional structures; the probability mass function of $\mu_{\alpha,x}^{(n)}$ depends smoothly on both $\alpha$ and $x$.
The purpose of this paper is to develop a fractional framework that encompasses both $\mu_{\alpha,x}^{(n)}$ and $B_{\alpha,n}$ in a unified manner based on the generalized binomial theorem. Although many of the resulting statements may appear elementary at first glance, their proofs often require techniques quite different from those used in the classical setting. These differences arise because, in general, $\mu_{\alpha,x}^{(n)}$ is not expressed as the law of the sum of i.i.d. random variables. Our main contribution is to clarify, in a systematic and unified way, the probabilistic and analytic features of $\mu_{\alpha,x}^{(n)}$ and $B_{\alpha,n}$. To the best of our knowledge, no previous work has constructed fractional analogues of the binomial distribution and the Bernstein operator in the unified way developed here.
1.2 Outline of the paper
The remainder of the paper is organized as follows.
The first part (Section 2). We establish explicit expressions and asymptotics for the moments of $\mu_{\alpha,x}^{(n)}$. In particular, we have
$\mathbb{E}[S_{\alpha,x}^{(n)}]=nx+O(e^{-\delta n})\quad\text{and}\quad\mathrm{Var}(S_{\alpha,x}^{(n)})=\frac{1}{\alpha}nx(1-x)+O(e^{-\delta n})$ (1.4)
as $n\to\infty$ for some $\delta>0$, where $S_{\alpha,x}^{(n)}$ is a random variable obeying $\mu_{\alpha,x}^{(n)}$. This reveals that the variance scales inversely with $\alpha$. See Figure 1. We also obtain a uniform bound of $Z_{\alpha,x}^{(n)}$ with respect to both $n\in\mathbb{N}$ and $x\in[0,1]$, and an explicit expression for the characteristic function of $\mu_{\alpha,x}^{(n)}$ with its quantitative estimate.
Figure 1: Variation of the fractional binomial distribution $\mu_{\alpha,x}^{(n)}$ with respect to $\alpha$. Shown are the probability mass functions for $n=20$ and $x=0.35$, where the parameter $\alpha$ takes the values $0.2$, $0.6$, 11, $1.4$, $1.8$, and $2.2$.
The second part (Section 3). Based on the properties given in Section 2, we prove several limit theorems: the weak law of large numbers, the central limit theorem, and the law of small numbers for $\mu_{\alpha,x}^{(n)}$. In particular, the limiting distribution in the latter case is a fractional variant of the Poisson distribution defined in terms of the Mittag-Leffler function (see e.g., [13, 6]).
The third part (Section 4). We investigate several properties of the fractional Bernstein operator $B_{\alpha,n}$ and its iterates. We prove the uniform approximation result and the Kelisky–Rivlin-type limit theorem for $B_{\alpha,n}$. We also prove the convergence of the appropriately scaled operators $(B_{\alpha,n}^{\lfloor nt\rfloor})_{n\in\mathbb{N}_{0}}$, $t\geq 0$, to the Wright–Fisher diffusion semigroup with generator $\mathcal{L}_{\alpha}f(x)=(2\alpha)^{-1}x(1-x)f^{\prime\prime}(x)$ on $C^{2}([0,1])$ as $n\to\infty$ in some sense. Furthermore, we extend this convergence to the case where the fractional order is a continuous function $\alpha(x)\colon[0,1]\to(0,\infty)$, and the limiting generator is given by $\mathcal{L}_{\alpha(\cdot)}f(x)=(2\alpha(x))^{-1}x(1-x)f^{\prime\prime}(x)$ on $C^{2}([0,1])$, which leads to more-general diffusion processes on $[0,1]$.
1.3 Future perspectives
We systematically generalize known results for the classical binomial distribution and Bernstein operator into their fractional counterparts by means of a unified framework based on the generalized binomial theorem. This approach enables us to make substantial contributions to the study of probability distributions and related fields. We expect that this framework will shed light on the analytical and geometric aspects of such structures, and deepen the understanding of various mathematical models, including stochastic processes, applied statistics, and information geometry. For example, the fractional binomial distributions may provide flexible models for count data which can describe both overdispersion and underdispersion, in view of (1.4). Furthermore, the study of the geometric properties of the statistical manifold formed by these distributions may lead to new developments in statistical theory and its applications.
1.4 Remarks
The terminology fractional in probability theory is used in several different contexts. For instance, a typical use can be seen in the notion of fractional Brownian motion. See [20] for the fractional binomial distribution in this sense. Yet another use is derived from fractional calculus. We refer to [5] for the fractional binomial process defined in terms of Caputo’s fractional derivatives. We also mention some existing notions of linear operators named the fractional Bernstein operator. In [24], the author studies the fractional Bernstein operator $(1-B_{n})^{\alpha}$, $\alpha>0$, which differs completely from ours. Another different formulation of the fractional Bernstein operator from ours can be found in [31]. Our fractional Bernstein operator $B_{\alpha,n}$ appears to be new, being directly derived from the generalized binomial theorem.
1.5 Notation